OpenStudy (anonymous):

how to exact value of sin(cos^-1(4/5)) ?

OpenStudy (anonymous):

im not sure what ur saying ??

hero (hero):

I'll see what I can do to help you

OpenStudy (anonymous):

ok...:)

OpenStudy (anonymous):

sr but i dont get where you got pi in "cosine 4pi/5" ?

OpenStudy (anonymous):

ok lets redo it, may be i copied the incorrect data

OpenStudy (anonymous):

lets inverse cosine 4/5=y so you have cos y= 4/5, find sin y, your original question was sin(inverse cosine 4/5)= sin (y)

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

thats the exact value ?

OpenStudy (anonymous):

you need to find sin 4/5, if you have a calculator

OpenStudy (anonymous):

im not allowed to use calculator

OpenStudy (anonymous):

is there anyway that i can find the value without calculator ?

OpenStudy (anonymous):

unfortunately not, well the closed we can it convert to radians by multiplying 4/5 with pi/180

hero (hero):

@psi9epsilon it's already in radians

OpenStudy (anonymous):

@hero, sorry i did not read it was in radians, well then that all about it, unless you want to convert it to degrees by multiplying with 180/pi

hero (hero):

Doing that won't help to find the exact value

OpenStudy (anonymous):

thats what I was trying to say, you cant get more out of sin (4/5) without a calculator

OpenStudy (anonymous):

@hero, do you disagree with the method, if so, kindly show an interesting solution , i am eager to learn : )

hero (hero):

Apparently you're supposed to use pythagorean trig identity to do it

zepdrix (zepdrix):

oo oh oh oh that makes sense :O

zepdrix (zepdrix):

 interesting c: